I AM TRYING TO HELP MY DAUGHTER SOLVE THE FOLLOWING PROBLEM:
x=3y
3y-6=2x
Can you help me to explain how to solve this for her? Where do I begin?
Thanks,
Jan

Found 2 solutions by venugopalramana, wuwei96815:
Answer by venugopalramana(3286)

x=3y...............I
3y-6=2x ............II
SUBSTITUTE FOR X FROM EQN.I IN TO EQN.II
3Y-6=2*(3Y)=6Y
3Y-6Y=6
-3Y=6
Y=6/-3=-2
SUBSTITUTING THIS IN EQN.I WE GET
X=3Y=3*(-2)=-6

Answer by wuwei96815(245)

There are several ways to solve this problem. Let's use the substitution method. Since we know the value of x (x=3y), we can substitute this value in the second equation and solve for y.

3y-6=2x
3y-6=2(3y)
3y-6=6y
-6=-3y+6y
-6=3y
-6/3=y
-2=y

Now that we have the value for y we can go back and substitute this in either equation and solve for x.
x=3y
x=3(-2)
x=-6

So, x=-6 and y=-2.